The beta-blocker drug, timolol, is expected to reduce the need for heart bypass surgery. Its
composition by mass is 49.4% C, 7.64% H, 17.7% N, 15.2% O, and 10.1% S. The mass of
0.0100 mol of timolol is 3.16 g. (a) What is the simplest formula of timolol? (b) What is
the molecular formula of timolol?
1
Expert's answer
2020-09-06T15:00:54-0400

a) C : H : N : O : S = 49.4/12 : 7.64/1 : 17.7/14 : 15.2/16 : 10.1/32 =

= 4.117 : 7.64 : 1.26 : 0.95 : 0.3156 =

= 13 : 24 : 4 : 3 : 1

the simplest formula of timolol is C13H24N4O3S "\implies" M( C13H24N4O3S) = 316 g/mol


b) M (molecul) = m / n = 3.16 / 0.0100 = 316 g/mol

M(simplest) = M (molecul) = 316 g/mol

It means that the molecular formula mathes to the simplest one: C13H24N4O3S